# Taxonomy of differentiationP. Komarov, F. van Breugel, J. N. Kutz

### v1

@@ 4.3.2 Recommendations
Finite Elements is not well suited to differentiating a typical array of numbers representing a signal, but it is the most versatile method for solving governing PDEs on arbitrary domains with arbitrary boundary conditions, which are common in engineering, especially in two and three dimensions. Due to the use of piecewise functions and the weak form, Finite Elements are also capable of representing true discontinuities like shockwaves, [^10] It is possible to approximate a discontinuity with Spectral methods, as demonstrated by Dedalus [dedalus], but the solution will require a high number of modes to represent and is still technically smooth, with rounded corners and a steep connecting segment. which can be present in the solutions of hyperbolic PDEs. Due to its complexity, Finite Elements is best done via established software. Commercial packages require a license, while the free, open-source FEniCSx requires formulating a PDE in the weak form [dokken].
@@ 5.1 Recommendations
When a signal is known to be periodic, applying an ideal low-pass filter, i.e. 2 with higher modes zeroed out before FFT-1, separates signal from high-frequency noise efficiently and optimally—although some noise spectrum will still unavoidably overlap the signal’s spectrum. Beware, one should not attempt to use the Chebyshev basis in the presence of noise, because polynomials have nonuniform oscillations and therefore do not filter noise equally well throughout the domain. Even worse, noise-induced fit inaccuracies propagate and compound from higher modes to lower ones during differentiation due to Chebyshev mode coupling [^11] By contrast with the aliasing of sinusoids, the best fit to samples from a single higher-degree polynomial requires many different lower-degree polynomials. [brown-cheb], resulting in systematic error toward the domain edges, as depicted in Figure 19. Likewise, the eigenmodes of a signal, ideal for parsimonious representation [ddse], are generally not best for separating noise due to their nonuniform frequency content across the domain. So in practice, with noisy, aperiodic data, one of the many methods discussed in the following two sections are likely to provide the most accurate results.
@@ 6.2 Bayesian Interpretation of the Kalman Filter
This form is also solvable with calculus. Taking the gradient w.r.t. $\mathbf{x}_{n}$ and setting equal to zero results in $\mathbf{\hat{x}}_{n}=(\mathbf{C}^{T}\mathbf{R}^{-1}\mathbf{C}+\mathbf{P}_{n|n-1}^{-1})^{-1}(\mathbf{C}^{T}\mathbf{R}^{-1}\mathbf{y}_{n}+\mathbf{P}_{n|n-1}^{-1}\mathbf{\hat{x}}_{n|n-1})$, which can be manipulated into $\mathbf{\hat{x}}_{n}=\mathbf{\hat{x}}_{n|n-1}+\mathbf{K}(\mathbf{y}_{n}-\mathbf{C}\mathbf{\hat{x}}_{n|n-1})$, although showing equivalence is tedious,141414unless the Woodbury matrix identity and clever factoring really light your fire see [johns-hopkins].
@@ 6.3 Forward-Backward Kalman Smoothing
where braces represent sets of all states, measurements, or inputs across steps, and $p_{\mathbf{w}},p_{\mathbf{v}}$ are the noise distributions. In this formulation, the prior of each state transition is conditioned on the raw previous state, not the previous state estimate as in the single-step case, because $\mathbf{x}_{n-1}$ is also being optimized over and is therefore available. Considering the joint probability distribution only up to $\mathbf{x}_{n}$ and marginalizing over all $\{\mathbf{x}_{0},...\mathbf{x}_{n-1}\}$ recovers Equation 24 [johns-hopkins].
@@
If the system dynamics, $\mathbf{A}$, are invertible, then it is also possible to run the Kalman filter, or whole RTS smoothing, on the data in reverse. Note $\mathbf{Q}$ is left as is for this procedure, because uncertainty due to entropy should be still be modeled as increasing across steps, even if time is reversed. Initial guesses $\mathbf{\hat{x}}_{0},\mathbf{P}_{0}$ bias Kalman estimates for several iterations, so combining forward and reverse passes can sometimes improve results at domain edges.
@@ 6.4 Generalized Kalman Filtering for Robust Estimation
Figure 23: RTS smoother (green) and robust MAP smoother (dashed pink) estimates on the cruise control data from Figure 22, but this time with 4 randomly chosen points turned into outliers. All matrices remain the same as before except for $\mathbf{Q}$, which has been scaled up by by a factor of $10^{5}$ to coerce the solution to deform significantly toward outliers, because RTS smoothing does remarkably well even with outliers when the system matrices are accurate. Despite such poor choice of $\mathbf{Q}$, a robust smoother with process penalty $\sqrt{2}\|\cdot\|_{1}$ and measurement penalty $1.04\sum\text{Huber}(\cdot,2)$ (constants explained in footnote) is able to ignore outliers and obtain an estimate close to ground truth.
@@ 6.4.1 Alternative Notion of Robustness
In controls and signal processing, the term “robust” is also used to mean guaranteed performance under input or model (parametric) uncertainty [robust-control, dullerud2000course, xie]. For unknown inputs, this gives rise to $H_{\infty}$ filtering/smoothing [nagpal, batteries], which is related to $H_{2}$ filtering [robust-control], of which the Kalman filter is the most famous example. [^16] To cast the Kalman filter into form of Equation 31, the salient output and internal state evolution are the error $\mathbf{e}$ and its dynamics. Disturbance can be split into pieces and multiplied by half-powers of covariance matrices $\mathbf{Q}$ and $\mathbf{R}$ to produce appropriately-shaped-but-deterministic (w.r.t. $\mathbf{\tilde{w}}$) process and measurement disturbances: $\begin{bmatrix}\mathbf{e}_{n}\\ \mathbf{e}_{n-1}\end{bmatrix}=\begin{bmatrix}(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{A}&\begin{bmatrix}(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{Q}^{1/2}&-\mathbf{K}\mathbf{R}^{1/2}\end{bmatrix}\\ \mathbb{I}&0\end{bmatrix}\begin{bmatrix}\mathbf{e}_{n-1}\\ \begin{bmatrix}\mathbf{\tilde{w}}_{\text{proc}}\\ \mathbf{\tilde{w}}_{\text{meas}}\end{bmatrix}_{n}\end{bmatrix}$ Using the time-domain version of the $H_{2}$ norm, which is defined with G’s impulse response (see [robust-control]), along with this ($\mathbf{\tilde{A}},\mathbf{\tilde{B}},\mathbf{\tilde{C}},\mathbf{\tilde{D}}$), we can manipulate the cost function of Equation 32 into $\text{Tr}(\mathbf{\tilde{C}}\mathbf{P}\mathbf{\tilde{C}}^{T})$, where $\mathbf{P}$ is the controllability Grammian $\sum_{m=0}^{\infty}\mathbf{\tilde{A}}^{m}\mathbf{\tilde{B}}\mathbf{\tilde{B}}^{T}(\mathbf{\tilde{A}}^{m})^{T}$, which solves the discrete-time Lyapunov equation, $\mathbf{\tilde{A}}\mathbf{P}_{n-1}\mathbf{\tilde{A}}^{T}-\mathbf{P}_{n}+\mathbf{\tilde{B}}\mathbf{\tilde{B}}^{T}=0$. Algebra recovers Equations 21 and 22, modulo the probabilistic bits. Choosing $\mathbf{\tilde{w}}\sim\mathcal{N}(0,\mathbb{I})$ puts the Grammian sum for $\mathbf{P}$ in the form of an expected value, $\mathbb{E}[\cdot]$, completing the bridge. These methods, perhaps confusingly, are unrelated to MAP generalizations of the Kalman filter. Instead, they consider a linear system named G, which maps exogenous “disturbance” $\mathbf{\tilde{w}}$ and internal state to output error $\mathbf{\tilde{z}}$ and next state (or state derivative in continuous time):
(31) $\begin{bmatrix}\mathbf{\tilde{x}}_{n+1}\\ \mathbf{\tilde{z}}_{n}\end{bmatrix}={\underbrace{\left[\begin{array}[]{c|c}\mathbf{\tilde{A}}&\mathbf{\tilde{B}}\\ \hline\cr\mathbf{\tilde{C}}&\mathbf{\tilde{D}}\end{array}\right]}_{\text{G}}}\begin{bmatrix}\mathbf{\tilde{x}}_{n}\\ \mathbf{\tilde{w}}_{n}\end{bmatrix}$
@@
However, this mathematically intense framework is often weaker than desired; $H_{p}$-optimal solutions are optimal regardless of any particular form of bounded-energy disturbance. This can be counterintuitive, especially because it treats disturbances as deterministic, whereas from the outset we have considered random noise (section 5). The classic Kalman filter is always $H_{2}$-optimal because it minimizes the system norm, but we understand it is also the “best” estimator in a probabilistic sense under certain conditions and have seen examples in section 6.4 of extending that same notion of best (MAP) to alternative assumptions, requiring filter modifications. Thus this $H_{2}$ sense of optimality may not actually be best for a given scenerio. Similarly, the $H_{\infty}$ filter is impressively robust but extremely conservative: Some systems that can be treated with Kalman filtering have unbounded $H_{\infty}$ norm, making $H_{\infty}$ filter design infeasible. Even when applicable, an $H_{\infty}$ filter places more weight on measurements and less on priors than a Kalman filter, allowing more high-frequency noise through [crassidis_junkins]. This notion of robustness is not the right tool for model-based filtering unless truly nothing is known about the disturbance, a rare situation. A smoother estimate, more appropriate for differentiation, will be possible under very bare assumptions, like wide uniform priors and anticipation of outliers.
@@ 7 Differentiating Noisy Data Without Prior Knowledge
In the absence of a model, estimating the derivative of a function, x(t)x(t), [^20] It is common in the context of noisy data to keep with the Kalman filter notation of section 6, where the symbol $y$ is reserved for noisy measurements and underlying signal is called $x$, so the independent variable must take on another symbol, $t$. This contrasts with earlier sections of this review, where functions are usually named $y$, taking input $x$. given noisy samples, $\mathbf{y}=\mathbf{x}+\boldsymbol{\eta}=x(\{t_{n}\})+\eta(\{t_{n}\})$, where $\{t_{n}\}$ is a set of sample points and $\eta$ is a noise process, is an ill-posed problem. In other words, the data are ambiguous, so it is not possible to solve for the derivative without imposing prior assumptions. Nevertheless, practitioners must solve this ill-posed problem routinely, often without much guidance. This has been increasingly true in data-driven science, where large-scale and often high-dimensional data streams are used to construct models [ddse].
@@ 7.1 Performance Metrics in the Presence and Absence of Ground Truth
A whole cornucopia of methods have been developed to estimate derivatives of noisy data, each with a set of tunable hyperparameters that affect its behavior, such as polynomial degree, window width, and number of iterations, which we collectively call $\Phi$. Differentiating in the absence of prior knowledge comes down to choosing a method and hyperparameter settings that achieve a good answer, where “good” entails two major concerns, each of which can be assigned a definite metric [floris2020]:
@@
    Accuracy: The derivative estimate, $\mathbf{\hat{\dot{x}}}$, should be faithful to the true derivative, $\mathbf{\dot{x}}$. We can measure this using the root mean squared error:
@@
Due to the complex relationship between $\Phi$ and $\mathbf{\hat{\dot{x}}}$ through a differentiation method, Equations 35 and 36 are unfortunately not convex in $\Phi$, nor are gradients with respect to $\Phi$ easy to find. But they can still be optimized via search strategies such as Nelder-Mead [nelder-mead], a downhill simplex direct search, starting from multiple initial conditions [pynumdiff, floris2020].
@@ 7.2 Prefiltering Methods
A common method of smoothing is to simply take a moving average, which is equivalent to convolving with a discrete kernel of uniform values that sum to 1. We can also take the median over windows, which can remove spurious values while preserving sharp transitions if noise levels are not too high, or we can convolve with alternative normalized kernels sampled from Gaussian or Friedrichs212121a bump function, e.g. $e^{-1/(1-x^{2})}$, $x\in(-1,1)$ functions. These kernel methods constitute crude digital low-pass filters, with frequency responses (magnitude Bode plots [oppenheim]) shown in the first three panels of Figure 27. The median filter is not visualized, because it is nonlinear and lacks a well-defined response.
@@
where $\omega$ is frequency, $\omega_{c}$ is the “cutoff” where power is halved, and $n$ is the filter order. This response, visualized in Figure 27, is intended to stay as flat as possible in the passband, thereby introducing minimal distortion, desirable because distortion can corrupt derivative estimates. In contrast to the global Fourier-Spectral Method (section 4.2.1), which can be used to perform ideal low-pass filtering by zeroing out higher modes prior to inverse transform, the Butterworth filter is local and can only achieve gradual cutoff. But as compensation for allowing attenuated high frequencies through, Butterworth filters are much more flexible, applicable to any signal, not only those that are periodic.

### v2

@@ 4.3.2 Recommendations
Finite Elements is not well suited to differentiating a typical array of numbers representing a signal, but it is the most versatile method for solving governing PDEs on arbitrary domains with arbitrary boundary conditions, which are common in engineering, especially in two and three dimensions. Due to the use of piecewise functions and the weak form, Finite Elements are also capable of representing true discontinuities like shockwaves, [^11] It is possible to approximate a discontinuity with Spectral Methods, as demonstrated by Dedalus [dedalus], but the solution will require a high number of modes to represent and is still technically smooth, with rounded corners and a steep connecting segment. which can be present in the solutions of hyperbolic PDEs [olver]. Due to its complexity, Finite Elements is best done via established software. Commercial packages require a license, while the free, open-source FEniCSx requires formulating a PDE in the weak form [dokken].
@@ 5.1 Recommendations
When a signal is known to be periodic, applying an ideal low-pass filter, i.e. 2 with higher modes zeroed out before FFT-1, separates signal from high-frequency noise efficiently and optimally—although some noise spectrum will still unavoidably overlap the signal’s spectrum. Beware, one should not attempt to use the Chebyshev basis in the presence of noise, because polynomials have nonuniform oscillations and therefore do not filter noise equally well throughout the domain. Even worse, noise-induced fit inaccuracies propagate and compound from higher modes to lower ones during differentiation due to Chebyshev mode coupling [^12] By contrast with the aliasing of sinusoids, the best fit to samples from a single higher-degree polynomial requires many different lower-degree polynomials. [brown-cheb], resulting in systematic error toward the domain edges, as depicted in Figure 19. Likewise, the eigenmodes of a signal, ideal for parsimonious representation [ddse], are generally not best for separating noise due to their nonuniform frequency content across the domain. So in practice, with noisy, aperiodic data, one of the many methods discussed in the following two sections is likely to provide the most accurate results.
@@ 6.2 Bayesian Interpretation of the Kalman Filter
This form is also solvable with calculus. Taking the gradient w.r.t. $\mathbf{x}_{n}$ and setting equal to zero results in $\mathbf{\hat{x}}_{n}=(\mathbf{C}^{T}\mathbf{R}^{-1}\mathbf{C}+\mathbf{P}_{n|n-1}^{-1})^{-1}(\mathbf{C}^{T}\mathbf{R}^{-1}\mathbf{y}_{n}+\mathbf{P}_{n|n-1}^{-1}\mathbf{\hat{x}}_{n|n-1})$, which can be manipulated into $\mathbf{\hat{x}}_{n}=\mathbf{\hat{x}}_{n|n-1}+\mathbf{K}(\mathbf{y}_{n}-\mathbf{C}\mathbf{\hat{x}}_{n|n-1})$, although showing equivalence is tedious,151515unless the Woodbury matrix identity and clever factoring really light your fire see [johns-hopkins].
@@ 6.3 Forward-Backward Kalman Smoothing
where braces represent sets of all states, measurements, or inputs across steps, and $p_{\mathbf{w}},p_{\mathbf{v}}$ are the noise distributions. In this formulation, the prior of each state transition is conditioned on the raw previous state, not the previous state estimate as in the single-step case, because $\mathbf{x}_{n-1}$ is also being optimized over and is therefore available. Considering the joint probability distribution only up to $\mathbf{x}_{n}$ and marginalizing over all $\{\mathbf{x}_{0},...,\mathbf{x}_{n-1}\}$ recovers Equation 24 [johns-hopkins].
@@
If the system dynamics, $\mathbf{A}$, are invertible, then it is also possible to run the Kalman filter, or whole RTS smoothing, on the data in reverse. Note $\mathbf{Q}$ is left as is for this procedure, because uncertainty due to entropy should still be modeled as increasing across steps, even if time is reversed. Initial guesses $\mathbf{\hat{x}}_{0},\mathbf{P}_{0}$ bias Kalman estimates for several iterations, so combining forward and reverse passes can sometimes improve results at domain edges.
@@ 6.4 Generalized Kalman Filtering for Robust Estimation
Figure 23: RTS smoother (green) and robust MAP smoother (dashed pink) estimates on the cruise control data from Figure 22, but this time with 4 randomly chosen points turned into outliers. All matrices remain the same as before except for $\mathbf{Q}$, which has been scaled up by a factor of $10^{5}$ to coerce the solution to deform significantly toward outliers, because RTS smoothing does remarkably well even with outliers when the system matrices are accurate. Despite such a poor choice of $\mathbf{Q}$, a robust smoother with process penalty $\sqrt{2}\|\cdot\|_{1}$ and measurement penalty $1.04\sum\text{Huber}(\cdot,2)$ (constants explained in footnote) is able to ignore outliers and obtain an estimate close to ground truth.
@@ 6.4.1 Alternative Notion of Robustness
In controls and signal processing, the term “robust” is also used to mean guaranteed performance under input or model (parametric) uncertainty [robust-control, dullerud2000course, xie]. For unknown inputs, this gives rise to $H_{\infty}$ filtering/smoothing [nagpal, batteries], which is related to $H_{2}$ filtering [robust-control], of which the Kalman filter is the most famous special case. [^17] To cast the Kalman filter into form of Equation 31, the salient output and internal state are both the error $\mathbf{e}$, with dynamics as derived in Equation 19. Vector-valued disturbance can be split into pieces and multiplied by half-powers of covariance matrices $\mathbf{Q}$ and $\mathbf{R}$ to produce appropriately-shaped-but-deterministic (w.r.t. $\mathbf{\tilde{w}}$) process and measurement disturbances: $\begin{bmatrix}\mathbf{e}_{n}\\ \mathbf{e}_{n-1}\end{bmatrix}=\begin{bmatrix}(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{A}&\begin{bmatrix}(\mathbb{I}-\mathbf{K}\mathbf{C})\mathbf{Q}^{1/2}&-\mathbf{K}\mathbf{R}^{1/2}\end{bmatrix}\\ \mathbb{I}&0\end{bmatrix}\begin{bmatrix}\mathbf{e}_{n-1}\\ \begin{bmatrix}\mathbf{\tilde{w}}_{\text{proc}}\\ \mathbf{\tilde{w}}_{\text{meas}}\end{bmatrix}_{n}\end{bmatrix}$ Using the time-domain version of the $H_{2}$ norm, which is defined with G’s impulse response (see [robust-control]), along with this ($\mathbf{\tilde{A}},\mathbf{\tilde{B}},\mathbf{\tilde{C}},\mathbf{\tilde{D}}$), we can manipulate the cost function of Equation 32 into $\text{Tr}(\mathbf{\tilde{C}}\mathbf{P}\mathbf{\tilde{C}}^{T})$, where $\mathbf{P}$ is the controllability Gramian $\sum_{m=0}^{\infty}\mathbf{\tilde{A}}^{m}\mathbf{\tilde{B}}\mathbf{\tilde{B}}^{T}(\mathbf{\tilde{A}}^{m})^{T}$, which solves the discrete-time Lyapunov equation, $\mathbf{\tilde{A}}\mathbf{P}_{n-1}\mathbf{\tilde{A}}^{T}-\mathbf{P}_{n}+\mathbf{\tilde{B}}\mathbf{\tilde{B}}^{T}=0$. Algebra recovers Equations 21 and 22, modulo the probabilistic bits. Choosing $\mathbf{\tilde{w}}\sim\mathcal{N}(0,\mathbb{I})$ puts the Gramian sum for $\mathbf{P}$ in the form of an expected value, $\mathbb{E}[\cdot]$, completing the bridge. These methods, perhaps confusingly, are unrelated to MAP generalizations of the Kalman filter. Instead, they consider a linear system named G, which maps exogenous “disturbance” $\mathbf{\tilde{w}}$ and internal state to output error $\mathbf{\tilde{z}}$ and next state (or state derivative in continuous time):
(31) $\begin{bmatrix}\mathbf{\tilde{x}}_{n+1}\\ \mathbf{\tilde{z}}_{n}\end{bmatrix}={\underbrace{\left[\begin{array}[]{c|c}\mathbf{\tilde{A}}&\mathbf{\tilde{B}}\\ \hline\cr\mathbf{\tilde{C}}&\mathbf{\tilde{D}}\end{array}\right]}_{\text{\normalsize G}}}\begin{bmatrix}\mathbf{\tilde{x}}_{n}\\ \mathbf{\tilde{w}}_{n}\end{bmatrix}$
@@
However, this mathematically intense framework is often weaker than desired: $H_{p}$-optimal solutions are optimal regardless of any particular form of bounded-energy disturbance. This can be counterintuitive, especially because it treats disturbances as deterministic, whereas from the outset we have considered random noise (section 5). The classic Kalman filter is always $H_{2}$-optimal because it minimizes the system norm, but we understand it is also the “best” estimator in a probabilistic sense under certain conditions and have seen examples in section 6.4 of extending that same notion of best (MAP) to alternative assumptions, requiring filter modifications. Thus this $H_{2}$ sense of optimality may not actually be best for a given scenario. Similarly, the $H_{\infty}$ filter is impressively robust but extremely conservative. Some systems that can be treated with Kalman filtering have unbounded $H_{\infty}$ norm, making $H_{\infty}$ filter design infeasible. Even when applicable, an $H_{\infty}$ filter places more weight on measurements and less on priors than a Kalman filter, allowing more high-frequency noise through [crassidis_junkins]. This notion of robustness is not the right tool for model-based filtering unless truly nothing is known about the disturbance, a rare situation. A smoother estimate, more appropriate for differentiation, will be possible under very bare assumptions, like wide uniform priors and anticipation of outliers.
@@ 7 Differentiating Noisy Data Without Prior Knowledge
In the absence of a model, estimating the derivative of a function, x(t)x(t), [^21] It is common in the context of noisy data to keep with the Kalman filter notation of section 6, where the symbol $y$ is reserved for noisy measurements and underlying signal is called $x$, so the independent variable must take on another symbol, $t$. This contrasts with earlier sections of this review, where functions are usually named $y$, taking input $x$. given noisy samples, $\mathbf{y}=\mathbf{x}+\boldsymbol{\eta}=x(\{t_{n}\})+\eta(\{t_{n}\})$, where $\{t_{n}\}$ is a set of sample points and $\eta$ is a noise process, is an ill-posed problem. In other words, the data are ambiguous, so it is not possible to solve for the derivative without imposing prior assumptions. Nevertheless, practitioners must solve this ill-posed problem routinely, often without much guidance. This has been increasingly true in data-driven science, where large-scale and often high-dimensional data streams are used to construct models [ddse, apriori-boulder].
@@ 7.1 Performance Metrics in the Presence and Absence of Ground Truth
A whole cornucopia of methods have been developed to estimate derivatives of noisy data, each with a set of tunable hyperparameters that affect its behavior, such as polynomial degree, window width, and number of iterations, which we collectively call $\Phi$. Differentiating in the absence of prior knowledge comes down to choosing a method and hyperparameter settings that achieve a good answer, where “good” entails two major concerns: closeness of fit and complexity [apriori-boulder]. Following [floris2020], we use the following metrics:
@@
    Accuracy: The derivative estimate, $\mathbf{\hat{\dot{x}}}$, should be faithful to the true derivative, $\mathbf{\dot{x}}$, resulting in a small root mean squared error:
@@
Due to the complex relationship between $\Phi$ and $\mathbf{\hat{\dot{x}}}$ through a differentiation method, Equations 35 and 36 are unfortunately not convex in $\Phi$, nor are gradients with respect to $\Phi$ easy to find. But they can still be optimized via strategies such as Nelder-Mead [nelder-mead], a downhill simplex direct search, starting from multiple initial conditions [pynumdiff, floris2020].
@@ 7.2 Prefiltering Methods
A common method of smoothing is to simply take a moving average, which is equivalent to convolving with a discrete kernel of uniform values that sum to 1. We can also take the median over windows, which can remove spurious values while preserving sharp transitions if noise levels are not too high, or we can convolve with alternative normalized kernels sampled from Gaussian or Friedrichs222222a bump function, e.g. $e^{-1/(1-x^{2})}$, $x\in(-1,1)$ functions. These kernel methods constitute crude digital low-pass filters, with frequency responses (magnitude Bode plots [oppenheim]) shown in the first three panels of Figure 27. The median filter is not visualized, because it is nonlinear and lacks a well-defined response.
@@
where $\omega$ is frequency, $\omega_{c}$ is the “cutoff” where power is halved, and $n$ is the filter order. This response, visualized in Figure 27, is intended to stay as flat as possible in the passband, thereby introducing minimal distortion. In contrast to the global Fourier-Spectral Method (section 4.2.1), which can be used as an ideal low-pass filter by zeroing out higher modes followed by inverse transforming, the Butterworth filter is local and can only achieve gradual cutoff. But as compensation, Butterworth filters are applicable to any signal, not only those that are periodic.